OpenStudy (anonymous):

A right cylinder has a radius r of 10.9 cm and a height h of 43.7 cm. What is the volume of the cylinder in m3?

OpenStudy (anonymous):

what is m3??

OpenStudy (anonymous):

\[piR^2h=m^3(v)\]

OpenStudy (anonymous):

I mean \[\=cm^2\] cubic centimeters

OpenStudy (anonymous):

\[\cm^3\]

OpenStudy (anonymous):

its vol is area of base times height.. pi*r^2*h..

OpenStudy (anonymous):

\[\pi(10.9)^2*43.7\]